OpenStudy (anonymous):

Solve: 4 over 5 y + 2 =6 A. 5 B. 7 C. 9 D. 11 I got 11, im not sure if that's right.

OpenStudy (anonymous):

If you are talking about (4/5)y + 2 = 6, you would solve it like this: (4/5)y + 2 = 6 (4/5)y = 4 y = 5

OpenStudy (anonymous):

well, there wasn't parenthesis in the equation.. maybe that's why i got it wrong. thanks.

OpenStudy (anonymous):

your welcome